Assignment 3 (Work, Energy and
Momentum) – Due Wednesday 20 May 2020
Question 1 (9 marks):
A block of mass 2.5 kg is pushed 2.2 m along a frictionless horizontal table by a
constant 16 N force directed 25° downward to the horizontal. Determine the work
done on the block by:
a) The applied force (3)
b) The normal force exerted by the table, and (2)
c) The gravitational force (2)
d) Determine the total work done (2)
25°
2.5 kg
1
Expert's answer
2020-05-15T08:38:02-0400

Given data

Mass of block m= 2.5 kg

Acceleration due to gravity g= 9.80 m/s2

Applied force is F= 16 N

----------------------------------------------------------------

a)

The work done by the applied force is

"W_F=Fd cos25^o=(16 N)(2.2 m) cos25=31.9 N= 32 J"

b)

The work done by the normal force exerted by the table, is

"W_N= Nd cos90^o=Nd(0) = 0 J"

c)

The work done by the gravitational force is

"W_G= F_Gd cos90^o=F_Gd = 0 J"

d)

The total work done is

"W_Total= W_F+W_N+ W_G= 32 J"